3. Suppose a consumer’s demand function for milk is x1 = 1/4 * m/p1, or (m/4p1), where p1 is the price of milk and x1 is the quantity of milk. Let the consumer’s income be $120/week and the price of milk be $3/quart. Suppose the price of milk falls to? $2/quart.

(a) (1) How much is the Slutsky compensation? (Specify whether it is given or taken away.)

(b) (2) How much is the Slutsky substitution effect?

(c) (2) How much is the Slutsky income effect?
Answer #1

③ Demand for milk x, = 1 * on de = Api 1 x 120 and M. = 10 quarts, 4 3 consumers income = m =120 Price P, = $3 /quart Demandincome a slutsky compensation is $40 taken toon consumers © Stutzky substitution ebfect. We have to find m* ie income necessa© income effect. AX, W124 m) – ulpm). = 212,120) – 341 2,110) -4.120 - + 10 = 10 = 1125